1. A ball is thrown straight up and passes point at a height of 40 m above its starting point in 5 s. What was its initial speed?
Answer #1

Kinematic equation


y=vi*t-(.5)gt^2


40 = vi(5) - (0.5)(9.8)(5)^2

40 = vi(5) - 122.5

vi(5) = 162.5


vi = 32.5 m/s